OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 7,28,112,448,...

OpenStudy (anonymous):

Multiply the term with 4 to get the next term.

OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 24, 36, 54, 81, … an=24(1.5)n − 1 an=24(0.5)n − 1 an=24(12)n − 1 an=24(18)n − 1